a jar contain 4 red and 6 white balls. 3 balls are drawn at random. in how many ways can at least one red ball be drawn?
please show the work.. i really need to understand the process of it

Answer by stanbon(75887)

a jar contain 4 red and 6 white balls. 3 balls are drawn at random. in how many ways can at least one red ball be drawn?
----
P(at least one red) = 1 - P(no red )
---
= 1 - 6C3/10C3
---
= 1 - 1/6
---
= 5/6
==========
Cheers,
Stan H.